OpenStudy (anonymous):

Jose is paying for a six dollar meal using bills in his wallet. He has four one dollar bills, three five dollar bills, and two ten dollar bills. If he selects two bills at random, one at a time from his wallet, what is the probability that he will choose a one dollar bill and a five dollar bill to pay for the meal? Show your work.

OpenStudy (triciaal):

4 -$1 + 3- $5 + 2 -$10 one method probability of $1 first = 4/9 probability of $5 next = 3/8 when combined, multiply so (4/9)*(3/8 ) =1/6 probability of $5 first = 3/9 probability of $1 next = 4/8 combined = (3/9)*(4/8)= 1/6 another using permutations and combinations don't remember clearly first you have 9 items (4 + 3 + 2) and probability of a $1 or $5 is 7/9 the probability for the 2nd draw is 6/8 =3/4

OpenStudy (anonymous):

thank you